Session: Jun 2013 (session 2)
Wavelength: 18 cm


To give feedback on this experiment, click here

Experiment Feedback

Hartebeesthoek (Jonathan Quick at Mon Jun 03 12:40:00 2013) Minor failures: FORM
Main Hh clock jumped on Mark5B during scan 4, resynced prior to scan 5. No other known problems.

Svetloe (Isaenko at Mon Jun 03 12:42:13 2013) Success:

Zelenchukskaya (Kiselev P. at Mon Jun 03 12:42:30 2013) Success:

Badary (Zhukov at Mon Jun 03 12:42:47 2013) Success:

Noto (C. Nocita at Mon Jun 03 12:55:14 2013) Success:

Medicina (A. Maccaferri at Mon Jun 03 13:08:25 2013) Success:

Urumqi (Mingfei Gao at Mon Jun 03 13:23:33 2013) Success:
The Tsys overflow due to noise source have not be inject. At present,the problem have been fixed.

Westerbork (Geert Kuper at Mon Jun 03 13:39:50 2013) Success:

Onsala (Michael Lindqvist at Tue Jun 04 09:47:37 2013) Success:
RFI, otherwise no known problems.

Effelsberg (Uwe Bach at Tue Jun 04 14:25:52 2013) Failure: ANT
Elevation control failure. Was solved during observations, but too late to start observing.

Torun (K. Borkowski at Wed Jun 05 05:36:27 2013) Minor failures:
Swapped polarizations seen in the first ftp-ed scan; amended before the second ftp correlation.

Urumqi (Peng Li at Fri Jun 07 08:45:06 2013) Minor failures:
Because MK5 failures, from the fifth scan begins normal observation.

Urumqi (Peng Li at Fri Jun 07 10:23:21 2013) Success: